OpenStudy (anonymous):

You plant a fruit Tree. The probability that the tree will survive its first winter is 0.8. If it survives the first winter, the probability that the tree will have fruit by its fourth year is 0.6. What is the probability that the tree will bear fruit by its fourth year??

OpenStudy (anonymous):

(show all work)

OpenStudy (anonymous):

i think if u add 0.8 and 0.6 and divide by 2 u will get the answer but i am not sure

OpenStudy (anonymous):

\[\frac{ 0.8+0.6 }{ 2 }\]

OpenStudy (anonymous):

i need someone who's sure of the answer :s

OpenStudy (anonymous):

1.4/2 7

OpenStudy (anonymous):

sorry 0.7

OpenStudy (anonymous):

Ignore that.

OpenStudy (anonymous):

So the concept you have to use here is that if I have two events, A and B, the probability that they will both happen, P(A&B), is P(A)*P(B). So for example, the probability to flip a heads and then roll a 6 on a die is P(H)*P(6) = (1/2)(1/6) = (1/12)

OpenStudy (anonymous):

i never learned that before- can you help me with that??

OpenStudy (anonymous):

Gladly. Do you understand what I'm saying?

OpenStudy (anonymous):

kind of, how do you have 2 events?

OpenStudy (anonymous):

There are a lot of ways you could have 2 events.

OpenStudy (anonymous):

question is not sufficient acc to what asked

OpenStudy (anonymous):

I want one thing to happen, and I want a second thing to also happen. Or I want one thing to happen and then after that, something else to happen.

OpenStudy (anonymous):

ok but in my math problem how would you set that up, i don't know which two events to plug into the equation

OpenStudy (anonymous):

Don't think of it as plugging into an equation. Slow down, understand the concept I'm explaining to you.

OpenStudy (anonymous):

sorry i really didn't understand this question @SmoothMath and @erica123

OpenStudy (anonymous):

If I have two events, and I want to know the probability that they will BOTH happen, I simply multiply the probability that either of them will happen. What's the probability that I will flip heads twice? The first event is that I'll flip heads, and the second event is that I'll flip heads again. What's P(H), the probability of flipping heads once? It's (1/2), or .5 So, what's the probability that I will flip heads twice? .5*.5 = .25

OpenStudy (anonymous):

@Bhagyashree that's okay, buddy.

OpenStudy (anonymous):

ok so once flipping the coin 1/2 and again if you flip it twice its another 1/2 so its 0.4 i understand that concept in this problem but in my problem its confusing :s

OpenStudy (anonymous):

oops 0.25!

OpenStudy (anonymous):

Well, not .4 It would be (1/4) or .25 Right =)

OpenStudy (anonymous):

yes my mistake i typed it wrong

OpenStudy (anonymous):

So, if you understand that concept, then you can apply it to this problem.

OpenStudy (anonymous):

but then there's two different numbers, 0.8 and 0.6 and it says in 4 years i dont know how to do that one

OpenStudy (anonymous):

We want to know the probability that the tree will bear fruit in its 4th year, but to do that, it has to survive its first year. So what is the probability it will survive and then ALSO bear fruit?

OpenStudy (anonymous):

0.6?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

i dont know how to do it?

OpenStudy (anonymous):

Yes you do. 2 events. What's the probability that they will both happen?

OpenStudy (anonymous):

It's exactly the same as the examples I gave you.

OpenStudy (anonymous):

i dont know how to do it with this one do i multiply the 0.6 or the 0.8?

OpenStudy (anonymous):

... Your question makes no sense.

OpenStudy (anonymous):

sorry :( which one do i use to figure out which ones will happen?

OpenStudy (anonymous):

like do i use the probability if it survives the first winter (0.8) or in 4 years (0.6)

OpenStudy (anonymous):

Erica, let me just give you like... a bunch of simple examples and see if you can catch on. The probability of flipping heads and then flipping heads is .5 * .5 = .25 The probability of rolling a 1 and then flipping heads is (1/6) * (1/2) = (1/12) The probability of flipping a heads and then rolling an even number is (1/2) * (3/6) = (3/12) = .25 The probability of pulling the ace of spades and then rolling a 3 is (1/52) * (1/6) = (1/312)

OpenStudy (anonymous):

Imagine that the probability of someone speeding being pulled over is .3, and the probability of someone who's been pulled over getting a ticket is .7. If I speed, the probability of me being pulled over and THEN getting a ticket is .3 * .7 = .21

OpenStudy (anonymous):

ok so the probability of the tree surviving and bearing fruit is 0.6*0.8= .48?

OpenStudy (anonymous):

im not sure if thats right

OpenStudy (anonymous):

Imagine that the probability of me forgetting to shower each day is .14. If I don't shower, the probability of someone noticing is .42. The probability of me forgetting to shower and THEN have someone notice is .14 * .42 = .00588

OpenStudy (anonymous):

That's right. It has to survive its first winter and THEN bear fruit. .8*.6

OpenStudy (anonymous):

ok i think i sorta got it now wait so that the answer? or is there more to it

OpenStudy (anonymous):

That's it. Super simple.

OpenStudy (anonymous):

If you have 2 events, and you want to know the probability of BOTH of them happening, just multiply 'em up!

OpenStudy (anonymous):

ok thanks so much! um i really dont want to ask you this but i have 3 others like this but are somewhat different can you please help me on that?

OpenStudy (anonymous):

Surely. Make sure you try on your own first.

OpenStudy (anonymous):

ok sure so do i just post them here or make a new question thread?

OpenStudy (anonymous):

Traditionally, posting a new question is the thang.

OpenStudy (anonymous):

ok